OpenStudy (anonymous):

Lt (1/n)^(1/ln n)=? n->infty

OpenStudy (anonymous):

first suppose C=lt (1/n)^(1/ln n) Taking ln both sides ln C=(1/ln n) ln(1/n) ln C=(1/ln n)(ln n^-1) ln C=(1/ln n)(-1 ln n) Now both ln n will cancel out result in -1 ln C=-1 Taking anti log both sides C=e^-1